OpenStudy (anonymous):

Which set of coefficients of the terms in the expansion of the binomial (x + y)3 is correct? 1, 3, 3, 1 1, 4, 4, 1 1, 2, 2, 1 1, 5, 5, 1

OpenStudy (campbell_st):

well if you split it out a bit its \[(x+ y)^3 = (x + y)^2 \times (x + y) =(x^2 + 2xy + y^2)(x + y)\] distribute that and't you'll find the answer..

OpenStudy (campbell_st):

another way is the 1st and last terms have a coefficient of 1 the 2nd and 2nd last terms have a coefficient that is the same as the power

OpenStudy (anonymous):

I see, so it would be 1,3,3,1

OpenStudy (campbell_st):

correct